# Q-Probe: Scaling Image Quality Assessment to High Resolution via Context-Aware Agentic Probing

### v1

@@ Abstract
Machine Learning, ICML, IQA, Agentic
[…]
figureChallenges in detecting subtle distortions via global perception versus local zooming. (a) Existing MLLMs fail to capture subtle local artifacts. (b) Even when visible via cropping, Semantic Robustness Bias causes models to ignore defects in key semantic areas (e.g., face). (c-d) Naive zooming leads to Logic Collapse, where the model misinterprets natural bokeh as blur (c) or falsely learns that Zooming implies Low Quality (d). (e) Q-Probe mimics human active viewing (Global Perception $\rightarrow$ Local Scrutiny $\rightarrow$ Critical Thinking) to correctly distinguish artifacts from natural effects. (f-g) Distribution of the Vista-Bench dataset and performance comparison showing Q-Probe’s superiority.
@@ 1 Introduction
Nevertheless, RL-based IQA methods suffer from a heavy reliance on global macroscopic views, making it difficult to precisely capture subtle local degradation features (Fig. Q-Probe: Scaling Image Quality Assessment to High Resolution via Context-Aware Agentic Probing(a)). Even when such artifacts are detected, they are frequently overlooked due to insufficient attentional allocation. In real-world scenarios, degradation in high-value semantic regions (e.g., faces or license plates) should trigger a severe penalty in the quality metric, whereas existing models frequently overestimate quality in such cases due to semantic robustness bias (Fig. Q-Probe: Scaling Image Quality Assessment to High Resolution via Context-Aware Agentic Probing(b)). Moreover, these methods are typically restricted to processing global quality. In high-resolution contexts, this coarse-grained approach fails to resolve small objects and subtle local artifacts, leading to marked biases in quality prediction.
@@
The recent emergence of the “Thinking with Images” (OpenAI, 2025) paradigm has spurred the use of multi-scale tools for high-resolution tiny object perception (Zheng et al., 2025; Wang et al., 2025a). While this offers a promising avenue for fine-grained perception in IQA, directly transposing this local zoom-in strategy presents severe challenges. On one hand, traditional visual zoom-in methods relying on region-level supervision benefit from the diversity of target objects, merely requiring SFT trajectories to encompass the target regions. In the IQA context, however, training exclusively on degraded regions is prone to inducing overfitting, causing the model to establish a spurious causal correlation that “cropping implies low quality” (Fig. Q-Probe: Scaling Image Quality Assessment to High Resolution via Context-Aware Agentic Probing(d)). Consequently, the model loses its discriminative capability regarding content, exhibiting a systematic low-score bias across all local views. On the other hand, although existing MLLMs excel at local semantic analysis, they often overlook the critical dependency of quality assessment on global context. Taking the Depth-of-Field (DoF) effect in photography as an example, a sharp foreground contrasted with a blurred background typically signifies high-quality imaging. Yet, when isolated from the holistic view and focused solely on the background, models tend to misinterpret natural optical bokeh as degradation (e.g., blur), thereby erroneously assigning a low quality score (Fig. Q-Probe: Scaling Image Quality Assessment to High Resolution via Context-Aware Agentic Probing(c)). Furthermore, the current IQA landscape lacks a benchmark dedicated to local fine-grained degradations in high-resolution scenarios, which significantly constrains the evolution of algorithms toward more refined perception.
@@
To address the aforementioned challenges, we propose Q-Probe (Fig. Q-Probe: Scaling Image Quality Assessment to High Resolution via Context-Aware Agentic Probing(e)), the first agentic IQA model grounded in the Thinking with Images paradigm and designed specifically for high-resolution scenarios. To enable such fine-grained perception given the lack of dedicated benchmarks, we first construct Vista-Bench (Fig. Q-Probe: Scaling Image Quality Assessment to High Resolution via Context-Aware Agentic Probing(f)), covering a wide range of scenes and local distortion patterns. To ensure the authenticity and precision of the synthesized artifacts, we employ the wavelet transform (Zhang, 2019) to decouple image structure from texture, selectively injecting degradations into texture-rich regions to simulate realistic impairments. We then leverage the advanced capabilities of Gemini-2.5 Pro (Comanici et al., 2025) to achieve hierarchical scores, calculating weighted assessments based on local degradation severity and the semantic significance of the regions. Finally, a rigorous human quality review and filtering process is conducted to guarantee the reliability of the benchmark.
@@ 3.3 Stage 1: Perception Alignment via Pre-RL
where $y_{ij}\in\{0,0.5,1\}$ represents the ground truth preference label derived from human ratings. This reward is designed to capture fine-grained distinctions in quality ranking. To ensure training stability, we minimize the GRPO objective $\mathcal{J}(\theta)$, which integrates the clipped surrogate advantage and a KL-divergence penalty:
$\mathcal{J}(\theta)$ $=-\mathbb{E}\Bigg[\min\left(\frac{\pi_{\theta}}{\pi_{old}}A_{k},\text{clip}\left(\frac{\pi_{\theta}}{\pi_{old}},1-\epsilon,1+\epsilon\right)A_{k}\right)$
@@ 3.5 Stage 3: Precision Pursuit via Decoupled Post-RL
[…]
@@
[…]
@@ 5 Conclusion
Appendix A Data Construction and Automated Annotation Details
To efficiently construct Vista-Bench and scale Image Quality Assessment (IQA) to high-resolution scenarios, we developed a high-throughput, asynchronous data processing pipeline. This system leverages Gemini-2.5-Pro to generate fine-grained, hierarchical quality annotations. This section details the processing workflow and the specific prompt engineering strategies employed.
A.1 Wavelet-Based Degradation Injection
Existing IQA datasets often rely on global distortions (e.g., Gaussian blur applied uniformly), which allow models to bypass genuine visual scrutiny by learning simple global statistical deviations. To prevent such “shortcut learning” and simulate realistic high-resolution defects, we employ a targeted degradation injection strategy based on the Discrete Wavelet Transform (DWT).
Structure-Texture Decoupling.
We utilize the Haar wavelet to decompose high-resolution source images into four distinct frequency subbands: the low-frequency approximation ($LL$) and three high-frequency detail components ($LH$, HLHL, HHHH). The $LL$ band preserves the global semantic structure and flat regions, while the high-frequency bands capture fine-grained textures and edges where high-resolution artifacts typically manifest, as shown in Figure 4.
Targeted Artifact Synthesis.
Instead of corrupting the entire image, we selectively inject degradations (e.g., compression ringing, sensor noise, and slight defocus) specifically into the high-frequency subbands. This strategy ensures that artifacts are embedded within complex texture regions—such as hair, foliage, or architectural details—making them visually subtle yet perceptually critical. By keeping the $LL$ band intact, we maintain the overall structural integrity of the scene while introducing realistic local impairments.
Reconstruction.
The final degraded samples are generated via the Inverse Discrete Wavelet Transform (IDWT). This process forces the IQA agent to actively “zoom in” and scrutinize local details to detect quality drops, effectively simulating the fine-grained perception required for 4K scenarios.
A.2 Asynchronous Annotation Pipeline
Given the high-resolution nature of the source images and the need for dense cropping, a sequential processing approach would be computationally prohibitive. We implemented a Python-based asynchronous framework tailored for high-concurrency interaction with the Multimodal LLM (MLLM). The pipeline consists of three critical stages:
  1. 1.
    Dynamic Image Preprocessing and Encoding: To optimize the balance between visual fidelity and token consumption, we implemented an adaptive resizing strategy. Original high-resolution images are resized such that their maximum edge does not exceed 1,024 pixels ($H_{max}=1024$) using Lanczos resampling. This ensures that global artifacts are preserved while significantly reducing bandwidth usage. Images are then serialized into Base64 format for API transmission.
  2. 2.
    Semaphore-Based Concurrency Control: To maximize throughput without exceeding API rate limits, the system utilizes an asynchronous semaphore mechanism (set to a concurrency limit of 5). This allows for parallel processing of image batches while maintaining system stability. The pipeline utilizes aiohttp for non-blocking network I/O, ensuring that image encoding and API waiting times do not bottle-neck the annotation process.
  3. 3.
    Robust Output Parsing: Since MLLM outputs can vary in formatting, we implemented a resilient parsing logic that utilizes regular expressions to extract JSON structures from mixed-text responses. This ensures valid data extraction even if the model wraps outputs in Markdown fences or custom XML-like tags.
A.3 Context-Aware Prompt Design
A critical challenge in Agentic IQA is preventing the model from hallucinating artifacts in clean, high-quality crops (e.g., natural bokeh). To address this, we designed a Two-Step Context-Aware Prompt that enforces a “Zero-Penalty” rule.
The prompt explicitly instructs the agent to first establish a global baseline score and then apply penalties only if specific, man-made degradations are detected. The structured system instructions provided to the agent are presented below.
Core Prompt for Context-Aware Annotation SYSTEM INSTRUCTION:
You are an advanced Image Quality Assessment Expert.
USER INSTRUCTION:
You are provided with multiple images.
Image 1 is the FULL ORIGINAL IMAGE. The remaining images are CROPPED VIEWS (zoomed-in details) from Image 1. CRITICAL KNOWLEDGE: TARGET DEGRADATIONS
You are specifically looking for 5 types of MAN-MADE degradations:
1. Mosaic: Pixelation blocks. 2. JPEG: Heavy compression ringing/blocks. 3. Posterize: Banding or flat color areas. 4. Scanlines: Horizontal black lines. 5. Grid: White/colored grid lines overlaid. All these 5 degradations must be artificial blocks, not natural textures. Step 1: Analyze Global Quality (Baseline)
Evaluate global tech-fidelity. Do NOT penalize natural bokeh/blur. Start with a Base Score (1.0 - 5.0). High-quality usually starts at 4.5+.
Step 2: Crop Analysis & “Zero-Penalty” Rule
Examine each crop carefully.
If a crop is Clean (no artificial degradation): Penalty = $0.$ If a crop is Degraded: Apply penalty based on location and severity. Output Format
Please strictly follow this JSON structure:
  "step1_base_score": 4.8,
  "step2_crop_details": [
      "crop_id": "Crop 1",
      "status": "Clean",
      "reasoning": "Natural leaf texture...",
      "penalty": 0.0
      "crop_id": "Crop 2",
      "status": "Degraded",
      "reasoning": "Distinct 8x8 mosaic blocks...",
      "penalty": 0.5
  "step3_summary": "Base score 4.8 minus penalties...",
  "final_score": 4.3
This hierarchical prompting strategy ensures that the final score (final_score) is mathematically derived from the baseline (step1_base_score) minus the localized penalties, effectively decoupling artistic blur from technical degradation.
Appendix B Inference Pipeline Implementation Details
To evaluate Q-Probe’s capability to detect fine-grained degradations in high-resolution images, we developed a specialized inference pipeline based on the vLLM framework. This pipeline supports multi-turn visual reasoning and dynamic tool invocation, allowing the model to actively ”zoom in” on suspicious regions. The implementation details are described below.
Figure 4: We utilize the Wavelet Transform to extract high-frequency components, serving as spatial guidance to concentrate the degradation injection within semantically rich foreground regions.
B.1 Engine Configuration and Resolution Management
We deployed the model using the vLLM engine to maximize inference throughput and memory efficiency. To accommodate the ultra-high resolution of Vista-Bench images (often exceeding 4K resolution), we customized the multi-modal processor configuration. Specifically, we set the min_pixels to $256\times 28\times 28$ and max_pixels to $2560\times 28\times 28$. This configuration ensures that the model preserves sufficient high-frequency details during the initial global encoding phase, preventing the loss of small artifacts (e.g., subtle mosaic blocks or grid lines). The inference was conducted using bfloat16 precision with a tensor parallel size of 2, distributed across NVIDIA RTX 4090 GPUs. We also enabled enable_prefix_caching to accelerate the processing of shared system prompts across multiple turns.
B.2 The Agentic Reasoning Loop
Unlike traditional static inference, our pipeline implements a dynamic, multi-turn ”Observe-Reason-Act” loop. The process is defined as follows:
  1. 1.
    Initialization: The session begins with the system prompt defining the five target degradation types (Mosaic, JPEG, Posterize, Scanlines, Grid) and the ”Zero-Penalty” rule. The global view of the high-resolution image is fed into the model.
  2. 2.
    Global Analysis and Hypothesis Generation: The model first evaluates the global technical fidelity. If it detects potential local degradations that are ambiguous in the global view, it generates a special XML tag <tool_call> containing the normalized bounding box coordinates for a closer inspection.
  3. 3.
    Dynamic Tool Execution: A regular expression parser monitors the model’s output stream. Upon detecting a crop_image_normalized function call, the pipeline pauses text generation and executes the cropping logic:
    $I_{crop}=\text{Crop}(I_{global},[x_{1},y_{1},x_{2},y_{2}])$ (9)
    The cropped region is extracted from the original high-resolution source (not the resized tokens) to ensure maximum clarity.
  4. 4.
    Contextual Update: The cropped image is encoded into Base64 format and appended to the conversation history as a new user message. The model then resumes generation, analyzing the new visual evidence to confirm or dismiss the degradation hypothesis.
  5. 5.
    Termination: This loop continues for a maximum of $T=6$ turns. The process terminates early if the model ceases to call tools or provides a final score.
B.3 State Management and Reproducibility
[…]
B.4 Prompt Engineering
To guide the agent in performing rigorous technical quality assessment, we designed a structured System Prompt. This prompt serves four key functions: Role Assignment, Defect Taxonomy, Workflow Enforcement, and Output Standardization.
The exact system prompt used in our inference pipeline is shown below.
System Prompt for Agentic IQA Inference You are an expert Image Quality Assessment (IQA) assistant. Your task is to analyze the image for technical degradations, specifically focusing on digital artifacts like: 1. Mosaic/Pixelation: Blocky structures. 2. JPEG Compression: Ringing artifacts and blocking. 3. Posterization: Color banding or loss of gradient. 4. Scanlines/Grid: Artificial line overlays. 5. Noise/Blur: General degradation (common in KADID10K). Inference Steps: 1. Global Analysis: Briefly describe the overall content and perceived quality. Is it sharp? Is it noisy? 2. Defect Hunting: Actively look for the artifacts listed above. 3. Zoom/Crop Strategy: If you suspect an artifact in a specific area (e.g., background textures, edges), use the crop_image_normalized tool to inspect it closely. 4. Final Scoring: Based on the severity of the defects found: 5.0: Excellent, no visible defects. 4.0: Good, minor defects visible only upon close inspection. 3.0: Fair, defects are annoying but content is visible. 2.0: Poor, defects significantly affect quality. 1.0: Bad, extremely degraded. CRITICAL OUTPUT RULE: At the very end of your final response, you MUST output the final score enclosed in double brackets, strictly like this: [[3.45]].
Appendix C Supervised Fine-Tuning Data Construction
To equip the Q-Probe agent with both robust visual perception and reasoning capabilities, we constructed a high-quality Supervised Fine-Tuning (SFT) dataset, named Probe-CoT-3K. This dataset is designed to teach the model not only to detect degradations but also to articulate its reasoning process in a structured manner. The construction process involves a rigorous template-based generation strategy that pairs diverse images with multi-turn conversation trajectories.
C.1 Data Format and Templating
Each data entry in Probe-CoT-3K follows a standardized multi-turn conversation format, simulating the interaction between a user and an expert IQA agent. The data structure is defined as follows:
  • System Prompt: A fixed instruction set that defines the agent’s persona, the specific degradation taxonomy (Mosaic, JPEG, Posterize, Scanlines, Grid), and the critical “Zero-Penalty” rule for clean crops.
  • Global Analysis Turn: The first turn involves the user presenting the full image. The assistant’s response includes a global quality assessment, a tentative base score, and—crucially—the invocation of the crop_image_normalized tool to inspect suspicious regions.
  • Local Scrutiny Turn: The subsequent turn simulates the system returning the cropped images. The assistant then analyzes these local views, identifies specific artifacts (if any), applies the appropriate penalties, and calculates the final quality score.
C.2 Trajectory Diversity and Quality Control
A key challenge in SFT for agentic tasks is avoiding “shortcut learning,” where the model blindly predicts low scores whenever a crop tool is used. To mitigate this, our dataset includes a balanced mix of three trajectory types:
  1. 1.
    Positive Trajectories: The agent zooms in on complex textures (e.g., foliage, fabric) and correctly identifies them as clean, applying zero penalty.
  2. 2.
    Negative Trajectories: The agent detects actual degradations (e.g., JPEG blocks) in the crops and applies penalties.
  3. 3.
    Global-Only Trajectories: For obviously low-resolution or low-quality images, the agent decides not to zoom in and provides a score directly based on the global view.
Table C.2 illustrates the template structures used for generating these reasoning trajectories, showcasing the distinction between the agentic zoom-in path and the direct scoring path.
Table 5: The template structures for the SFT data. To ensure model robustness, we construct two distinct reasoning paths: (a) The Agentic Zoom-in Path for high-resolution images requiring local scrutiny, and (b) The Direct Scoring Path for low-resolution.
(a) High-Res / Agentic Path (With Crop Tool Usage)
[…]
Appendix D More Details about Vista-Bench
The significant performance superiority of Q-Probe over existing global-view IQA methods on Vista-Bench can be attributed to the fundamental paradigm shift from passive observation to active Context-Aware Agentic Probing. While traditional methods rely on coarse-grained global views that often miss high-frequency details, our framework introduces a nuanced evaluation capability centered on two key advantages:
1. Fine-Grained Perception vs. Coarse Global Scoring: Existing MLLMs typically process images from global perspective, causing subtle degradations in high-resolution scenarios (e.g., $4K$ inputs) to vanish during encoding. These models effectively ”guess” the quality based on a low-fidelity global impression. In contrast, Q-Probe utilizes an agentic zooming mechanism to actively localize and inspect microscopic degradations. Crucially, our method does not merely detect defects; it evaluates the perceptual impact of the degradation relative to the semantic importance of the region. By integrating the severity of local defects with the saliency of the content (e.g., applying severe penalties for artifacts on a foreground subject while adhering to a ”Zero-Penalty” rule for natural background bokeh), Q-Probe aligns its scoring logic with human visual attention, a capability largely absent in global-view baselines.
2. Generalization to Unseen Degradations: To rigorously validate that the performance gains stem from genuine visual understanding rather than distribution overfitting, we strategically incorporated unseen degradation types into the Vista-Bench test set—distortion patterns that the model was never exposed to during training. Standard baselines often fail on such out-of-distribution data as they rely on memorizing specific artifact features. However, Q-Probe demonstrates sustained high performance on these unseen samples. This proves that our model has transcended simple pattern matching and has acquired a robust, abstract capability to assess how generic visual anomalies impair the overall integrity of an image, regardless of whether the specific degradation pattern has been previously encountered.
Appendix E Limitations and Bad Cases Study
To comprehensively evaluate the boundaries of Q-Probe, we analyze both successful applications and failure modes. As demonstrated in Figure 5, under standard high-resolution scenarios, our agent exhibits superior performance. The context-aware cropping strategy successfully distinguishes between artificial degradations and natural depth-of-field effects (bokeh). By accurately localizing low-quality regions while ignoring non-subject blurry backgrounds, the model adheres strictly to the ”Zero-Penalty” rule, ensuring fair and accurate quality assessment.
Figure 5: Success Case: The agent effectively identifies the main subject and applies the “Zero-Penalty” rule to the natural bokeh in the background, demonstrating the effectiveness of our context-aware cropping strategy in standard scenarios.
However, robustness challenges arise in scenes exhibiting complex semantic dependencies. The first limitation involves semantic entanglement between foreground and background. As illustrated in Figure 6, while the model ignores standalone blurry backgrounds, it struggles when distant backgrounds contain textures strongly correlated with the subject’s narrative (e.g., a complex street scene behind a vehicle).
In these cases, the agent’s attention mechanism becomes confused by the high correlation. Instead of focusing solely on the subject, it erroneously interprets complex background textures as ”artificial noise” or ”processing artifacts.” This leads to erroneous cropping on irrelevant areas and subsequently applies unjustified penalties, indicating a need for a more sophisticated semantic understanding of scene composition.
Figure 6: Failure Mode I (Semantic Entanglement): While the agent can distinguish bokeh, its discriminative capability weakens when the background is strongly correlated with the subject (e.g., a car on a street). The model may misinterpret complex background textures as artifacts, leading to erroneous cropping.
The second limitation pertains to the insensitivity to spatial context differences, particularly between indoor and outdoor environments. As shown in Figure 7, our agent is prone to uniformly assigning minimal penalties to degradations in all background regions. This generalization is flawed because indoor backgrounds cannot be treated equivalently to vast outdoor settings.
Due to the confined spatial nature of indoor scenes, background artifacts remain perceptually significant compared to distant outdoor backgrounds. Consequently, while the penalty for indoor background degradation should not be as severe as foreground defects, it must not be negligible either. Our current model fails to apply this necessary moderate penalty, often underestimating the impact of background noise in spatially constrained environments.
Figure 7: Failure Mode II (Spatial Context Insensitivity): The agent incorrectly applies minimal penalties to indoor background degradations. Unlike vast outdoor settings, indoor scenes are spatially confined, meaning background artifacts remain perceptually significant and require a moderate penalty rather than a negligible one.
Finally, a third limitation arises from ambiguity in subject definition, leading to a salience bias during crop selection. As depicted in Figure 8, the agent incorrectly identifies the distant house as the primary subject—likely due to its status as the sole distinct architectural element—while relegating the river to the background.
This classification ignores the fact that the river occupies nearly half the frame and dominates the visual experience. In the context of Image Quality Assessment (IQA), degradations on such a prominent surface significantly impair perceptual quality. By classifying the river as a non-salient background, the model generates a flawed reasoning trajectory and fails to penalize visible defects, resulting in an inflated quality score.
Figure 8: Failure Mode III (Salience Bias): Ambiguity in scene composition introduces a bias: the agent misidentifies the small distant house as the main subject due to object uniqueness, while dismissing the dominant river as background. Consequently, the model fails to penalize degradations on the water surface, which are actually critical to perceptual quality.

### v2

@@ Abstract
Machine Learning, ICML
Figure 1: Challenges in detecting subtle distortions via global perception versus local zooming. (a) Existing MLLMs fail to capture subtle local artifacts. (b) Even when visible via cropping, Semantic Robustness Bias causes models to ignore defects in key semantic areas (e.g., face). (c-d) Naive zooming leads to Logic Collapse, where the model misinterprets natural bokeh as blur (c) or falsely learns that Zooming implies Low Quality (d). (e) Q-Probe mimics human active viewing (Global Perception $\rightarrow$ Local Scrutiny $\rightarrow$ Critical Thinking) to correctly distinguish artifacts from natural effects. (f-g) Data distribution of the high-resolution Vista-Bench and performance comparison showing Q-Probe’s superiority.
@@ 1 Introduction
Nevertheless, RL-based IQA methods suffer from a heavy reliance on global macroscopic views, making it difficult to precisely capture subtle local degradation features (Fig. 1(a)). Even when such artifacts are detected, they are frequently overlooked due to insufficient attentional allocation. In real-world scenarios, degradation in high-value semantic regions (e.g., faces or license plates) should trigger a severe penalty in the quality metric, whereas existing models frequently overestimate quality in such cases due to semantic robustness bias (Fig. 1(b)). Moreover, these methods are typically restricted to processing global quality. In high-resolution contexts, this coarse-grained approach fails to resolve small objects and subtle local artifacts, leading to marked biases in quality prediction.
@@
The recent emergence of the “Thinking with Images” (OpenAI, 2025) paradigm has spurred the use of multi-scale tools for high-resolution tiny object perception (Zheng et al., 2025; Wang et al., 2025a). While this offers a promising avenue for fine-grained perception in IQA, directly transposing this local zoom-in strategy presents severe challenges. On one hand, traditional visual zoom-in methods relying on region-level supervision benefit from the diversity of target objects, merely requiring SFT trajectories to encompass the target regions. In the IQA context, however, training exclusively on degraded regions is prone to inducing overfitting, causing the model to establish a spurious causal correlation that “cropping implies low quality” (Fig. 1(d)). Consequently, the model loses its discriminative capability regarding content, exhibiting a systematic low-score bias across all local views. On the other hand, although existing MLLMs excel at local semantic analysis, they often overlook the critical dependency of quality assessment on global context. Taking the Depth-of-Field (DoF) effect in photography as an example, a sharp foreground contrasted with a blurred background typically signifies high-quality imaging. Yet, when isolated from the holistic view and focused solely on the background, models tend to misinterpret natural optical bokeh as degradation (e.g., blur), thereby erroneously assigning a low quality score (Fig. 1(c)). Furthermore, the current IQA landscape lacks a benchmark dedicated to local fine-grained degradations in high-resolution scenarios, which significantly constrains the evolution of algorithms toward more refined perception.
@@
To address the aforementioned challenges, we propose Q-Probe (Fig. 1(e)), the first agentic IQA model grounded in the Thinking with Images paradigm and designed specifically for high-resolution scenarios. To enable such fine-grained perception given the lack of dedicated benchmarks, we first construct Vista-Bench (Fig. 1(f)), covering a wide range of scenes and local distortion patterns. To ensure the authenticity and precision of the synthesized artifacts, we employ the wavelet transform (Zhang, 2019) to decouple image structure from texture, selectively injecting degradations into texture-rich regions to simulate realistic impairments. We then leverage the advanced capabilities of Gemini-2.5 Pro (Comanici et al., 2025) to achieve hierarchical scores, calculating weighted assessments based on local degradation severity and the semantic significance of the regions. Finally, a rigorous human quality review and filtering process is conducted to guarantee the reliability of the benchmark.
@@ 3.3 Stage 1: Perception Alignment via Pre-RL
where $y_{ij}\in\{0,0.5,1\}$ represents the ground truth preference label derived from human ratings. This reward is designed to capture fine-grained distinctions in quality ranking. To ensure training stability, we maximize the GRPO objective $\mathcal{J}(\theta)$, which integrates the clipped surrogate advantage and a KL-divergence penalty:
$\mathcal{J}(\theta)$ $=\mathbb{E}\Bigg[\min\left(\frac{\pi_{\theta}}{\pi_{old}}A_{k},\text{clip}\left(\frac{\pi_{\theta}}{\pi_{old}},1-\epsilon,1+\epsilon\right)A_{k}\right)$
@@ 3.5 Stage 3: Precision Pursuit via Decoupled Post-RL
BRISQUE (Mittal et al., 2012a) 0.152 0.614 0.429 0.242 0.548 0.385 0.497 0.409
NIQE (Mittal et al., 2012b) 0.187 0.676 0.487 0.357 0.532 0.421 0.533 0.456
MUSIQ (Ke et al., 2021) 0.295 0.720 0.647 0.317 0.670 0.473 0.494 0.516
UNIQUE (Zhang et al., 2021) 0.310 0.751 0.513 0.393 0.703 0.649 0.608 0.561
MANIQA (Yang et al., 2022) 0.325 0.745 0.760 0.338 0.589 0.213 0.422 0.484
Qwen2.5-VL-7B (Bai et al., 2025) 0.385 0.848 0.787 0.390 0.787 0.754 0.735 0.669
LIQE (Zhang et al., 2023) 0.342 0.815 0.809 0.371 0.718 0.684 0.653 0.627
DeQA-Score (You et al., 2025) 0.398 0.852 0.831 0.383 0.756 0.677 0.738 0.662
Q-Align (Wu et al., 2023) 0.360 0.767 0.832 0.406 0.769 0.573 0.682 0.627
UnifiedReward-T (Wang et al., 2025d) 0.412 0.871 0.841 0.399 0.788 0.820 0.722 0.693
Q-Insight (Li et al., 2025) 0.365 0.872 0.856 0.429 0.816 0.806 0.749 0.699
VisualQuality-R1 (Wu et al., 2025) 0.451 0.875 0.871 0.469 0.848 0.855 0.805 0.739
Q-Probe (Ours) 0.728 0.892 0.901 0.474 0.829 0.871 0.837 0.790
@@
BRISQUE (Mittal et al., 2012a) 0.165 0.624 0.451 0.259 0.546 0.400 0.541 0.426
NIQE (Mittal et al., 2012b) 0.192 0.683 0.415 0.314 0.516 0.439 0.560 0.445
MUSIQ (Ke et al., 2021) 0.288 0.666 0.622 0.347 0.695 0.435 0.434 0.498
UNIQUE (Zhang et al., 2021) 0.305 0.708 0.548 0.361 0.729 0.590 0.581 0.546
MANIQA (Yang et al., 2022) 0.318 0.753 0.780 0.364 0.617 0.257 0.448 0.505
Qwen2.5-VL-7B (Bai et al., 2025) 0.362 0.854 0.806 0.420 0.837 0.810 0.772 0.694
LIQE (Zhang et al., 2023) 0.335 0.814 0.817 0.347 0.748 0.652 0.653 0.623
DeQA-Score (You et al., 2025) 0.375 0.858 0.873 0.381 0.793 0.703 0.743 0.675
Q-Align (Wu et al., 2023) 0.348 0.779 0.862 0.381 0.794 0.612 0.694 0.638
UnifiedReward-T (Wang et al., 2025d) 0.395 0.846 0.877 0.421 0.873 0.804 0.745 0.708
Q-Insight (Li et al., 2025) 0.346 0.872 0.881 0.462 0.851 0.829 0.794 0.719
VisualQuality-R1 (Wu et al., 2025) 0.427 0.878 0.821 0.458 0.871 0.840 0.843 0.734
Q-Probe (Ours) 0.776 0.900 0.892 0.476 0.876 0.863 0.813 0.799